Store can purchase shirts for seven dollars each each shirt is sold for $18 if the store wants to make a profit of $748 how many shirts message so stable extra present state the equation and then state the answer
1
Expert's answer
2020-09-10T19:15:42-0400

Let the number of shirts be "x"

The equation for the Cost =profit + buying price (7x)

=748+7x

The equation for the the revenue =18x ,x being the number of shirts sold.


To get the break even point ;

We equate Cost to Revenue i.e Cost=Revenue

748+7x=18x

18x-7x=748

11x=748

x=68 shirts